# Universal 3-Edge-Weightability of Regular Graphs

### v1

@@ 2. Proof of Theorem 1
Let $E_{i+1}$ be the set of edges with weight 0 in the subgraph $G[\cup_{j\geq i+1}I_{j}]$. If $E_{i+1}\neq\emptyset$, we process the edges in $E_{i+1}$ one by one. When considering an edge $e$, if both endpoints are hungry, we set $w(e)=d_{2}$; otherwise we leave $e$ unweighted. After processing all edges in $E_{i+1}$, let $U_{i+1}$ be the set of hungry vertices in $\cup_{j\geq i+1}I_{j}$. By construction, if an edge has both endpoints in $U_{i+1}$, then it has weight $d_{2}$; hence the induced subgraph $G[U_{i+1}]$ (if nonempty) has all edges weighted with $d_{2}$.
Take an arbitrary vertex $u\in U_{i+1}$, and suppose $u\in I_{\alpha}$ for some $\alpha\in[i+1,\ell]$. Since $u$ is hungry, $d_{w}(u)=jd_{2}$ for some $j\leq\alpha-2$. By assumption (C), for each $\beta\in[i+1,\alpha-1]$, $u$ is incident to at least one edge of weight $d_{2}$ with other endpoint in $I_{\beta}$. There are $(\alpha-1)-(i+1)+1=\alpha-i-1$ such indices $\beta$. Thus at most $(\alpha-2)-(\alpha-i-1)=i-1$ edges of weight $d_{2}$ incident to $u$ can have their other endpoint in $\cup_{j\geq\alpha+1}I_{j}$. Recall that $G[U_{i+1}]$ has all edges weighted with $d_{2}$. So within the subgraph $G[U_{i+1}]$, the vertex $u$ has at most i1i-1 neighbors in $U_{i+1}\cap(\cup_{j\geq\alpha+1}I_{j})$. Since $I_{\alpha}$ is an independent set and $u$ is arbitrary, it follows that $G[U_{i+1}]$ is (i1)(i-1)-degenerate. By Lemma 1, $U_{i+1}$ can be partitioned into at most $i$ independent sets in $G[U_{i+1}]$. Each such independent set is also independent in $G$. Hence we can write $U_{i+1}=J_{1}\cup\cdots\cup J_{r}$ with $r\leq i$, where each $J_{t}$ is independent in $G$.
By Lemma 2, for each $t\in[1,r]$ there exists a matching $M^{\prime}_{i,t}$ in the bipartite graph $G[I_{i},J_{t}]$ that saturates $J_{t}$. For every edge $e$ in $\bigcup_{t=1}^{r}M^{\prime}_{i,t}$, set $w(e)=d_{2}$. At this point, each vertex $v\in I_{i}$ is incident to at most $i$ edges of weight $d_{2}$ from these matchings, so $d_{w}(v)=jd_{2}$ for some $j\leq r\leq i$.
@@
    every vertex in $I_{\alpha}$ for each $\alpha\geq i$ is Type I, Type II or hungry;
@@
    for $\alpha\geq i+1$, if $I_{\alpha}$ has hungry vertices, then: for every hungry vertex $u$ in $I_{\alpha}$ and every $\beta\in[i,\alpha-1]$, $u$ is incident to at least one edge of weight $d_{2}$ whose other endpoint lies in $I_{\beta}$.
@@
Proceeding in this way for $i=\ell-2,\ell-3,\ldots,1$, we obtain at the end of Phase 1 that every vertex in $\cup_{\alpha\geq 1}I_{\alpha}$ is Type I or II. Since at the end, on one hand, each $v$ in $I_{1}$ has weighted degree $d_{2}-d_{1}$ (Type II) or 0 (Type I). On the other hand, if some $u$ in $I_{\alpha}$ ($\alpha\geq 2$) is hungry, it should satisfies (C’): $u$ is incident to at least one edge of weight $d_{2}$ with other endpoint in $I_{\beta}$ for each $\beta\in[1,\alpha-1]$ (at the end $i=1$). That is, $d_{w}(u)\geq(\alpha-1)d_{2}$ and $u$ is not hungry, a contradiction. Consequently, condition (1’) holds and there is no conflict within the subgraph $G[\cup_{i\geq 1}I_{i}]$. Moreover, every vertex $v\in I_{0}$ satisfies $d_{w}(v)\leq 0$ at this stage and (2’) holds.
For (3’), if there exists a Type I vertex $v$ in $I_{1}$ that has a Type I neighbor $u$ in some $I_{i}$ ($i\geq 2$), let $e_{v}$ and $e_{u}$ be the edges joining $v$ and $u$, respectively, to $I_{0}$ (recall that each vertex in $\cup_{j\geq 1}I_{j}$ has a neighbor in $I_{0}$ since $I_{0}$ is a maximum independent set). Clearly, $e_{v},e_{u},uv$ have not been weighted since $v,u$ are Type I (moreover, $v$ is Type I and in $I_{1}$). Set $w(e_{v})=w(e_{u})=-d_{1}$ and set $w(uv)=d_{2}$. Then both $u,v$ become Type II (so condition (1’) remains valid), and every vertex in $I_{0}$ still has weighted degree at most 0 (so condition (2’) remains valid). Repeat this step until no Type I vertex in $I_{1}$ has a Type I neighbor in $\cup_{i\geq 2}I_{i}$, and (3’) holds.
@@
For $i=0,1$, let $I_{i,0}=\{v\in I_{i}:d_{w}(v)=0\}.$ After Phase 1, by (1’) and (2’), the only possible conflicts in $G$ are those belonging to the bipartite subgraph $G[I_{0,0},I_{1,0}]$. If $G[I_{0,0},I_{1,0}]$ is empty, we are done. Assume therefore that it is nonempty. For each vertex $x\in I_{1,0}$, choose a neighbor $y_{x}$ (if exists) of $x$ in $I_{0,0}$.
Remark on dynamics: During Phase 2, the sets $I_{1,0}$, $I_{0,0}$, and the classifications based on them will change as we assign new weights. We will carefully track these changes to ensure the algorithm terminates.
At this point, every vertex in $I_{1,0}$ has all its neighbors either in $I_{0}$ or in Type II vertices of $\cup_{i\geq 2}I_{i}$ by (1’) and (3’).
We classify the vertices of $I_{1,0}$ according to their number of neighbors in $I_{0}$:
$I_{1,0}=X_{2^{+}}\cup X_{1}=X_{3^{+}}\cup X_{2}\cup X_{1},$
where $X_{i^{+}}=\{x\in I_{1,0}:|N(x)\cap I_{0}|\geq i\}$ and $X_{i}=\{x\in I_{1,0}:|N(x)\cap I_{0}|=i\}$ for $i\geq 1$.
Now consider the bipartite subgraph $H^{\prime}:=G[X_{2}\cup X_{1},\cup_{i\geq 2}I_{i}]$. On one hand, for any $x\in X_{2}\cup X_{1}$:
  • If $x\in X_{2}$, then $|N(x)\cap I_{0}|=2$, so $x$ has at least k2k-2 neighbors in $\cup_{i\geq 2}I_{i}$ (since all neighbors are either in $I_{0}$ or Type II vertices of $\cup_{i\geq 2}I_{i}$).
  • If $x\in X_{1}$, then $|N(x)\cap I_{0}|=1$, so $x$ has at least $k-1\geq k-2$ neighbors in $\cup_{i\geq 2}I_{i}$.
For any Type II vertex $z\in\cup_{i\geq 2}I_{i}$, by definition it is incident to exactly one edge of weight $d_{1}$ (to a vertex in $I_{0}$) and exactly $i$ edges of weight $d_{2}$ for some $i\geq 2$. Since all edges incident to vertices in $I_{1,0}$ currently have weight 0, any edge of weight $d_{2}$ incident to $z$ must connect $z$ to a vertex not in $I_{1,0}$. Thus $z$ has at least one neighbor not in $I_{1,0}$ via an edge of weight $d_{2}$, plus the edge of weight $d_{1}$ to $I_{0}$. Consequently, $z$ can be adjacent to at most k2k-2 vertices in $I_{1,0}$ (so to at most k2k-2 vertices in $X_{2}\cup X_{1}$).
Thus, in $H^{\prime}$, every vertex in $X_{2}\cup X_{1}$ has degree at least k2k-2, and every vertex in $\cup_{i\geq 2}I_{i}$ has degree at most k2k-2. To verify Hall’s condition for a matching saturating $X_{2}\cup X_{1}$, take any subset $S\subseteq X_{2}\cup X_{1}$. The number of edges incident to $S$ in $H^{\prime}$ is at least |S|(k2)|S|(k-2). Since each vertex in the neighborhood $N_{H^{\prime}}(S)$ has degree at most k2k-2, we have
$|S|(k-2)\leq e(S,N_{H^{\prime}}(S))\leq|N_{H^{\prime}}(S)|(k-2).$
Dividing by k2>0k-2>0 (as $k\geq 3$) gives $|S|\leq|N_{H^{\prime}}(S)|$. Hence Hall’s condition holds, and there exists a matching $M$ in $H^{\prime}$ that saturates $X_{2}\cup X_{1}$. Write $M=M_{1}\cup M_{2}$, where $M_{i}$ saturates $X_{i}$ for $i=1,2$.
@@
Let $X^{\ast}_{2^{+}}\subseteq X_{2^{+}}$ be the set of vertices that have a neighbor in $I_{0,0}$. If $X^{\ast}_{2^{+}}=\emptyset$, then we are done.
We process vertices from $X_{2^{+}}$ iteratively. Each iteration removes one vertex from $X_{2^{+}}$ (and hence from $X^{\ast}_{2^{+}}$). Although $X^{\ast}_{2^{+}}$ may expand due to vertices re-entering $I_{0,0}$ in Case 1, the set $X_{2^{+}}$ itself strictly decreases in size with each iteration. Therefore, the loop terminates after at most $|X_{2^{+}}|$ iterations.
Let $i=1$. While $X^{\ast}_{2^{+}}\neq\emptyset$, we select a vertex $x_{i}$ from $X^{\ast}_{2^{+}}$ (arbitrarily for $i=1$, and from the updated $X^{\ast}_{2^{+}}$ for $i>1$) and handle it according to the cases below. After processing $x_{i}$, it is removed from $I_{1,0}$ (since its weighted degree becomes non-zero), and consequently from $X_{2^{+}}$ and $X^{\ast}_{2^{+}}$ (this process will gradually eliminates vertices from $X_{2^{+}}$ until $X^{\ast}_{2^{+}}$ becomes empty). We then update $I_{1,0}$, $X_{2^{+}}$, $X^{\ast}_{2^{+}}$ accordingly.
@@
    for each previously processed vertex $x_{j}$ ($j<i$), one of the following items holds:
    • (i’)
@@
    • (ii’)
      $d_{w}(x_{j})=-3d_{1}$ and every neighbor $y$ of $x_{j}$ in $I_{0}$ satisfies $d_{w}(y)\neq-3d_{1}$.
@@
We distinguish three cases.
Case 1: $x_{i}\in X_{2^{+}}$, and $x_{i}$ has a neighbor $y^{\ast}\in I_{0}$ with $d_{w}(y^{\ast})\in(-\infty,-d_{1})\setminus\{-d_{2}-3d_{1}\}$. Set $w(x_{i}y^{\ast})=d_{2}$ and $w(x_{i}y_{x_{i}})=-d_{1}$. After these assignments, $d_{w}(x_{i})=d_{2}-d_{1}$ (hence condition (iii)(i’) holds for $x_{i}$), $d_{w}(y_{x_{i}})=-d_{1}$, and the weighted degree of $y^{\ast}$ increases by $d_{2}$. Since before the assignment $d_{w}(y^{\ast})<-d_{1}$, after adding $d_{2}$ we have $d_{w}(y^{\ast})<d_{2}-d_{1}$, which fulfills condition (ii). Furthermore, because originally $d_{w}(y^{\ast})\neq-d_{2}-3d_{1}$, we have $d_{w}(y^{\ast})\neq-3d_{1}$. Thus $d_{w}(y^{\ast})\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$, and conditions (ii) and (iii) continue to hold for $i+1$.
Case 2: $x_{i}\in X_{3^{+}}$ and every neighbor of $x_{i}$ in $I_{0}$ has weighted degree in $d_{1},d_{2}-d_{1})\cup\{-d_{2}-3d_{1}\}$. Since $x_{i}\in X_{3^{+}}$, it has at least three neighbors in $I_{0}$. Arbitrarily choose three edges $x_{i}y_{1},x_{i}y_{2},x_{i}y_{3}$ in $G[I_{0},X_{3^{+}}]$ and set $w(x_{i}y_{1})=w(x_{i}y_{2})=w(x_{i}y_{3})=-d_{1}$. This gives $d_{w}(x_{i})=-3d_{1}$, while $d_{w}(y_{1}),d_{w}(y_{2}),d_{w}(y_{3})\in[-2d_{1},d_{2}-2d_{1})\cup\{-d_{2}-4d_{1}\}$. Then every neighbor of $x_{i}$ in $I_{0}$ has weighted degree in $2d_{1},d_{2}-d_{1})\cup\{-d_{2}-4d_{1},-d_{2}-3d_{1}\}$; this set is consistent with (ii), and does not include $3d_{1}$ and so (iii) also holds for $i+1$.
Case 3: $x_{i}\in X_{2}$ and every neighbor of $x_{i}$ in $I_{0}$ has weighted degree in $d_{1},d_{2}-d_{1})\cup\{-d_{2}-3d_{1}\}$. Let $y^{\ast}$ be the unique neighbor of $x_{i}$ in $I_{0}$ other than $y_{x_{i}}$. Let $x_{i}z$ be the edge in $M_{2}$ that saturates $x_{i}$, where $z\in I_{q}$ for some $q\geq 2$. By (1’) and (3’), $z$ is Type II and so $d_{w}(z)=qd_{2}-d_{1}$. Now set $w(x_{i}y_{x_{i}})=w(x_{i}y^{\ast})=w(x_{i}z)=-d_{1}$. This gives $d_{w}(x_{i})=-3d_{1}$, while $d_{w}(y_{x_{i}})=-d_{1}$, $d_{w}(z)=qd_{2}-2d_{1}$ (that is, $z$ becomes Type III and so condition (1) remains valid), and $d_{w}(y^{\ast})\in[-2d_{1},d_{2}-2d_{1})\cup\{-d_{2}-4d_{1}\}$. Then $\{d_{w}(y^{\ast}),d_{w}(y_{x_{i}})\}\subseteq[-2d_{1},d_{2}-d_{1})\cup\{-d_{2}-4d_{1}\}$; this set is consistent with (ii), and does not include $3d_{1}$ and so (iii) also holds for $i+1$.
@@
Remark on the dynamics of $I_{0,0}$: In Case 1, it is possible that $d_{w}(y^{\ast})$ becomes 0 after adding $d_{2}$ (if originally $d_{w}(y^{\ast})=-d_{2}$), placing $y^{\ast}$ back into $I_{0,0}$. However, by assumption (i), all edges incident to $\{x_{j}:j>i\}$ are still unweighted, so any new conflict involving $y^{\ast}$ and $x_{j}$ ($j>i$) will be resolved when $x_{j}$ is processed. Moreover, $y^{\ast}$ cannot conflict with already processed vertices $x_{p}$ ($p<i$) because those have weighted degrees either $d_{2}-d_{1}$ or $3d_{1}$, both different from 0. For cases 2 and 3, all modified weighted degrees in $I_{0}$ become negative; so $I_{0}$-vertices cannot re-enter $I_{0,0}$.
Remark on the dynamics of $X^{\ast}_{2^{+}}$: Since (in Case 1) $I_{0,0}$ may expand, $X^{\ast}_{2^{+}}$ may also expand. However, $X^{\ast}_{2^{+}}$ is included in $X_{2^{+}}$, and $X_{2^{+}}$ decreases at each step. So the recursive process will terminate with $X_{2^{+}}=\emptyset$.
@@
After handling $x_{i}$, update $I_{0,0},X_{2^{+}}$ and $X^{\ast}_{2^{+}}$ respectively. Note that $x_{i}$ and all previously processed vertices $x_{p}$ ($p<i$) are no longer in $X^{\ast}_{2^{+}}$ (no longer in $X_{2^{+}}$ either), since they either have weighted degree $d_{2}-d_{1}$ (which differs from 0), or have weighted degree $3d_{1}$ while all their neighbors in $I_{0}$ have weighted degrees different from $3d_{1}$. If $X^{\ast}_{2^{+}}=\emptyset$, we are done. If $X^{\ast}_{2^{+}}\neq\emptyset$ let $x_{i+1}\in X^{\ast}_{2^{+}}$ and the assumptions (i)–(iii) hold for $i+1$. Proceeding and update $X^{\ast}_{2^{+}}$ repeatedly, we eventually eliminate all conflicts in $G[I_{0},X_{2^{+}}]$. Moreover, (1), (2), (3) remain valid currently. Hence no conflict remains incident to $X_{2^{+}}$.

### v2

@@ 2. Proof of Theorem 1
Let $E_{i+1}$ be the set of edges with weight 0 in the subgraph $G[\cup_{j\geq i+1}I_{j}]$. If $E_{i+1}\neq\emptyset$, we process the edges in $E_{i+1}$ one by one. When considering an edge $e$, if both endpoints are hungry, we set $w(e)=d_{2}$; otherwise we leave $e$ unweighted. After processing all edges in $E_{i+1}$, let $U_{i+1}$ be the set of hungry vertices in $\cup_{j\geq i+1}I_{j}$. By construction, if an edge has both endpoints in $U_{i+1}$, then it has already been weighted with $d_{2}$; hence the induced subgraph $G[U_{i+1}]$ (if nonempty) has all edges weighted with $d_{2}$.
Take an arbitrary vertex $u\in U_{i+1}$, and suppose $u\in I_{\alpha}$ for some $\alpha=i+1,\ldots,\ell$. Since $u$ is hungry, $d_{w}(u)=jd_{2}$ for some $j\leq\alpha-2$. By assumption (C), for each $\beta=i+1,\ldots,\alpha-1$, $u$ is incident to at least one edge of weight $d_{2}$ with other endpoint in $I_{\beta}$. There are $(\alpha-1)-(i+1)+1=\alpha-i-1$ such indices $\beta$. Thus at most $(\alpha-2)-(\alpha-i-1)=i-1$ edges of weight $d_{2}$ incident to $u$ can have their other endpoints in $\cup_{j\geq\alpha+1}I_{j}$. Recall that $G[U_{i+1}]$ has all edges weighted with $d_{2}$. So within the subgraph $G[U_{i+1}]$, the vertex $u$ has at most i1i-1 neighbors in $U_{i+1}\cap(\cup_{j\geq\alpha+1}I_{j})$. Since $I_{\alpha}$ is an independent set and $u$ is arbitrary, it follows that $G[U_{i+1}]$ is (i1)(i-1)-degenerate. By Lemma 1, $U_{i+1}$ can be partitioned into at most $i$ independent sets in $G[U_{i+1}]$. Each such independent set is also independent in $G$. Hence we can write $U_{i+1}=J_{1}\oplus\cdots\oplus J_{r}$ with $r\leq i$, where each $J_{t}$ is independent in $G$.
By Lemma 2, for each $t=1,2,\ldots,r$, there exists a matching $M^{\prime}_{i,t}$ in the bipartite graph $G[I_{i},J_{t}]$ that saturates $J_{t}$. For every edge $e$ in $\bigcup_{t=1}^{r}M^{\prime}_{i,t}$, set $w(e)=d_{2}$. At this point, each vertex $v\in I_{i}$ is incident to at most $r$ edges of weight $d_{2}$ from these matchings, so $d_{w}(v)=jd_{2}$ for some $j\leq r\leq i$.
@@
    for each $\alpha\geq i$, every vertex in $I_{\alpha}$ is Type I, Type II or hungry;
@@
    for $\alpha\geq i+1$, if $I_{\alpha}$ has hungry vertices, then: for every hungry vertex $u$ in $I_{\alpha}$ and every $\beta=i,\ldots,\alpha-1$, $u$ is incident to at least one edge of weight $d_{2}$ whose other endpoint lies in $I_{\beta}$.
@@
Proceeding in this way for $i=\ell-2,\ell-3,\ldots,1$, we obtain at the end that every vertex in $\cup_{\alpha\geq 1}I_{\alpha}$ is Type I or II. Since at the end, on one hand, each $v$ in $I_{1}$ has weighted degree $d_{2}-d_{1}$ (Type II) or 0 (Type I). On the other hand, if some $u$ in $I_{\alpha}$ ($\alpha\geq 2$) is hungry, it should satisfies (C’): $u$ is incident to at least one edge of weight $d_{2}$ with other endpoint in $I_{\beta}$ for each $\beta=1,\ldots,\alpha-1$ (at the end $i=1$). That is, $d_{w}(u)\geq(\alpha-1)d_{2}$ which implies that $u$ is not hungry, a contradiction. Consequently, condition (1’) holds. Moreover, every $y$ in $I_{0}$ satisfies $d_{w}(y)\in(-\infty,0]$ at this stage and (2’) holds.
For (3’), if there exists a Type I vertex $v$ in $I_{1}$ that has a Type I neighbor $u$ in some $I_{i}$ ($i\geq 2$), let $e_{v}$ and $e_{u}$ be the edges joining $v$ and $u$ to $I_{0}$, respectively (recall that each vertex in $\cup_{j\geq 1}I_{j}$ has a neighbor in $I_{0}$ since $I_{0}$ is a maximum independent set). Clearly, $e_{v},e_{u}$ and $uv$ have not been weighted since $v$ and $u$ are both Type I (moreover, $v$ is in $I_{1}$). Set $w(e_{v})=w(e_{u})=-d_{1}$ and set $w(uv)=d_{2}$. Then both $u$ and $v$ become Type II (so condition (1’) remains valid), and every vertex in $I_{0}$ still has weighted degree at most 0 (so condition (2’) remains valid). Repeat this step until no Type I vertex in $I_{1}$ has a Type I neighbor in $\cup_{i\geq 2}I_{i}$, and (3’) holds finally.
@@
For $i=0,1$, let $I_{i,0}=\{v\in I_{i}:d_{w}(v)=0\}.$ By our construction, all edges incident to vertices in $I_{0,0}\cup I_{1,0}$ currently have weight 0. After Phase 1, by (1’) and (2’), the only possible conflicts in $G$ are those belonging to the bipartite subgraph $G[I_{0,0},I_{1,0}]$. If $G[I_{0,0},I_{1,0}]$ is empty, we are done. Assume therefore that it is nonempty.
For each $x\in I_{1,0}$ with $N(x)\cap I_{0,0}\neq\emptyset$, choose an arbitrary vertex from $N(x)\cap I_{0}$ and denote it by $y_{x}$. For $i\geq 1$, let $X_{i}=\{x\in I_{1,0}:|N(x)\cap I_{0}|=i\}$ and $X_{i^{+}}=\cup_{j\geq i}X_{j}$.
Now consider the bipartite subgraph $H^{\prime}:=G[X_{1}\cup X_{2},\cup_{i\geq 2}I_{i}]$. On one hand, for any $x\in X_{1}\cup X_{2}$, the vertex $x$ has at least k2k-2 Type II neighbors in $\bigcup_{i\geq 2}I_{i}$. This holds because, by conditions (1’) and (3’), every neighbor of $x$ either lies in $I_{0}$ or is a Type II vertex in $\bigcup_{i\geq 2}I_{i}$. On the other hand, consider a Type II vertex $z\in I_{i}$ with $i\geq 2$. By definition, $z$ is incident to exactly $i$ edges of weight $d_{2}$ and one edge of weight $d_{1}$. Since all edges incident to vertices in $X_{1}\cup X_{2}$ currently carry weight 0, any edge of weight $d_{2}$ or $d_{1}$ incident to $z$ must have its other endpoint outside $X_{1}\cup X_{2}$. Thus $z$ can be adjacent to at most k2k-2 vertices in $X_{1}\cup X_{2}$. Thus, in $H^{\prime}$, every vertex in $X_{1}\cup X_{2}$ has degree at least k2k-2, and every vertex in $\cup_{i\geq 2}I_{i}$ has degree at most k3k-3. By Hall’s theorem, there exists a matching $M$ in $H^{\prime}$ that saturates $X_{1}\cup X_{2}$. Write $M=M_{1}\oplus M_{2}$, where $M_{i}$ saturates $X_{i}$ for $i=1,2$. For each $x\in X_{1}\cup X_{2}$, let $z_{x}$ be the (Type II) vertex in $\cup_{i\geq 2}I_{i}$ such that $xz_{x}\in M$.
@@
Let $X^{\ast}_{2^{+}}\subseteq X_{2^{+}}$ be the set of vertices that have a neighbor in $I_{0,0}$. If $X^{\ast}_{2^{+}}=\emptyset$, then we are done. Suppose $X^{\ast}_{2^{+}}\neq\emptyset$.
We select a vertex $x_{i}$ from $X^{\ast}_{2^{+}}$ (arbitrarily for $i=1$, and from the updated $X^{\ast}_{2^{+}}$ for $i>1$) and handle it according to the cases below. After processing $x_{i}$, it will be removed from $I_{1,0}$ (since its weighted degree will become non-zero), and consequently from $X_{2^{+}}$ and $X^{\ast}_{2^{+}}$ (this process will gradually eliminates vertices from $X_{2^{+}}$ until $X^{\ast}_{2^{+}}$ becomes empty). We then update $I_{1,0}$, $X_{2^{+}}$, $X^{\ast}_{2^{+}}$ accordingly.
@@
    each previously processed vertex $x_{j}$ ($j<i$) satisfies one of the following conditions:
@@
    • $d_{w}(x_{j})=-3d_{1}$ and every neighbor $y$ of $x_{j}$ in $I_{0}$ satisfies $d_{w}(y)\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$.
@@
We distinguish two cases.
Case 1: If there exists $y^{\ast}\in N(x_{i})\cap I_{0}$ such that $d_{w}(y^{\ast})\in(-\infty,-d_{1})\setminus\{-d_{2}-3d_{1}\}$. Clearly, $y^{\ast}\neq y_{x_{i}}$, since by the definition of $y_{x_{i}}$, we have $d_{w}(y_{x_{i}})=0$ at this stage. Set $w(x_{i}y^{\ast})=d_{2}$ and $w(x_{i}y_{x_{i}})=-d_{1}$. After these assignments, $d_{w}(x_{i})=d_{2}-d_{1}$, while $d_{w}(y_{x_{i}})=-d_{1}$ and $d_{w}(y^{\ast})\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$. Moreover, the weighted degree of each vertex in $(I_{0}\setminus\{y_{x_{i}},y^{\ast}\})\cup(X_{2^{+}}\setminus\{x_{i}\})$ remains unchanged. Thus, conditions (i)-(iii) continue to hold for $i+1$.
Case 2: If every vertex in $N(x_{i})\cap I_{0}$ has weighted degree in $d_{1},d_{2}-d_{1})\cup\{-d_{2}-3d_{1}\}$.
Subcase 2.1: If $x_{i}\in X_{3^{+}}$. Let $y_{1},y_{2},y_{3}\in N(x_{i})\cap I_{0}$ and set $w(x_{i}y_{1})=w(x_{i}y_{2})=w(x_{i}y_{3})=-d_{1}$. This gives $d_{w}(x_{i})=-3d_{1}$, while $d_{w}(y_{1}),d_{w}(y_{2}),d_{w}(y_{3})\in[-2d_{1},d_{2}-2d_{1})\cup\{-d_{2}-4d_{1}\}$. Then every vertex in $N(x_{i})\cap I_{0}$ has weighted degree in $2d_{1},d_{2}-d_{1})\cup\{-d_{2}-4d_{1},-d_{2}-3d_{1}\}$, which is a subset of $(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$. Moreover, the weighted degree of each vertex in $(I_{0}\setminus\{y_{1},y_{2},y_{3}\})\cup(X_{2^{+}}\setminus\{x_{i}\})$ remains unchanged. Thus, conditions (i)-(iii) continue to hold for $i+1$.
Subcase 2.2: If $x_{i}\in X_{2}$. Assume $N(x_{i})\cap I_{0}=\{y_{x_{i}},y^{\prime}\}$ and $z_{x_{i}}\in I_{q}$ for some $q\geq 2$. Recall that $z_{x_{i}}$ is Type II and so $d_{w}(z_{x_{i}})=qd_{2}-d_{1}$. Now set $w(x_{i}y_{x_{i}})=w(x_{i}y^{\prime})=w(x_{i}z_{x_{i}})=-d_{1}$. This gives $d_{w}(x_{i})=-3d_{1}$, while $d_{w}(y_{x_{i}})=-d_{1}$, $d_{w}(y^{\prime})\in[-2d_{1},d_{2}-2d_{1})\cup\{-d_{2}-4d_{1}\}$ and $d_{w}(z_{x_{i}})=qd_{2}-2d_{1}$ (that is, $z_{x_{i}}$ becomes Type III and so condition (1) remains valid). Then $\{d_{w}(y_{x_{i}}),d_{w}(y^{\prime})\}\subseteq[-2d_{1},d_{2}-2d_{1})\cup\{-d_{2}-4d_{1}\}$, which is a subset of $(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$. Moreover, the weighted degree of each vertex in $(I_{0}\setminus\{y_{x_{i}},y^{\prime}\})\cup(X_{2^{+}}\setminus\{x_{i}\})$ remains unchanged. Thus, conditions (i)-(iii) continue to hold for $i+1$.
@@
Remark on the dynamics of $I_{0,0}$: In Case 1, it is possible that $d_{w}(y^{\ast})$ becomes 0 after adding $d_{2}$ (if originally $d_{w}(y^{\ast})=-d_{2}$), placing $y^{\ast}$ back into $I_{0,0}$. However, by assumption (i), all edges incident to $\{x_{j}:j>i\}$ are still unweighted, so any new conflict involving $y^{\ast}$ and $x_{j}$ ($j>i$) will be resolved when $x_{j}$ is processed. Moreover, $y^{\ast}$ cannot conflict with already processed vertices $x_{s}$ ($s<i$) because those have weighted degrees either $d_{2}-d_{1}$ or $3d_{1}$, both different from 0. For Case 2, all modified weighted degrees in $I_{0}$ become negative; so $I_{0}$-vertices cannot re-enter $I_{0,0}$.
@@
After handling $x_{i}$, update $I_{0,0},X_{2^{+}}$ and $X^{\ast}_{2^{+}}$, respectively. Note that $x_{i}$ and all previously processed vertices $x_{j}$ ($j<i$) are no longer in $X^{\ast}_{2^{+}}$ (no longer in $X_{2^{+}}$ either), since they either have weighted degree $d_{2}-d_{1}$ (which differs from 0), or have weighted degree $3d_{1}$ while all their neighbors in $I_{0}$ have weighted degrees different from $3d_{1}$. If $X^{\ast}_{2^{+}}=\emptyset$, we are done. If $X^{\ast}_{2^{+}}\neq\emptyset$ let $x_{i+1}\in X^{\ast}_{2^{+}}$ and the assumptions (i)-(iii) hold for $i+1$. Proceeding and update $X^{\ast}_{2^{+}}$ repeatedly, we eventually eliminate all conflicts in $G[I_{0},X_{2^{+}}]$. Moreover, conditions (1)-(3) remain valid currently. Hence no conflict remains incident to $X_{2^{+}}$ finally.
Remark on the dynamics of $X^{\ast}_{2^{+}}$: Since (in Case 1) $I_{0,0}$ may expand, $X^{\ast}_{2^{+}}$ may also expand. However, $X^{\ast}_{2^{+}}$ is included in $X_{2^{+}}$, and $X_{2^{+}}$ decreases at each step. So the recursive process will terminate with $X^{\ast}_{2^{+}}=\emptyset$.